OpenStudy (anonymous):

Two cars leave Denver at the same time and travel in opposite directions. One car travels 10mi/h faster than the other car. The cars are 300 mi apart in 3hrs. How fast is each car traveling?

OpenStudy (anonymous):

r t = d Two cars moving in opposite directions is the same as one car moving at their combined speed in one direction for 300 miles. Solve the following for x, the speed of the slowest car.\[3 (2 x+10)=300 \] x = 45 mph. x + 10 = 55 mph

OpenStudy (anonymous):

(45 + 55) 3 = 300